# Explicitly Modeling Subcortical Vision with a Neuro-Inspired Front-End Improves CNN Robustness

### v1

@@ 3.1 EVNets Improve Neuronal and Behavioral Alignment with Primate Vision
We evaluated whether coupling the SubcorticalBlock with the VOneBlock improved alignment with primate V1. As illustrated in Figure 2, the inclusion of the SubcorticalBlock upstream of the VOneBlock introduces hallmark extra-classical RF response properties absent from the VOneBlock alone. In particular, we observe an increased surround suppression in the size tuning curve and clear contrast saturation in the contrast-sensitivity curve.
Figure 2: Subcortical preprocessing improves explanability of extra-classical RF properties in V1. SF, size, and contrast tuning curves (left to right) for an example neuron in the VOneBlock with and without subcortical preprocessing (M cell). Example frames from the drifting gratings stimulus set are shown below. V1 cells in isolation exhibit predominantly classical response properties but when coupled with subcortical processing exhibit emergent behaviors consistent with those empirically observed, such as enhanced surround modulation and contrast saturation[doi:10.1152/jn.00692.2001, SCLAR19901].
Motivated by these observations, we used the BrainScore platform [SchrimpfKubilius2018BrainScore, Schrimpf2020integrative] to quantitatively evaluate V1 alignment. As shown in Table 1, EVResNet50 achieves higher scores on nearly all response properties, besides improving both the mean response property score and V1 neural predictivity scores. Notably, the highest gains are observed for the surround modulation and RF size tuning — both associated with the increased surround suppression. A tradeoff in SF tuning and texture modulation is observed, which can be attributed to the fact that no changes were done to the GFB to account for the upstream processing.
Table 1: EVResNet50 overperforms baselines on V1 neural predictivity, V1 response property alignment, and shape bias. BrainScore [SchrimpfKubilius2018BrainScore, Schrimpf2020integrative] V1 alignment scores and shape bias [geirhos2022imagenettrained] for ResNet50, VOneResNet50, and EVResNet50. Values indicate mean $\pm$ SD ($n=3$ seeds).
@@
ResNet50 [he2015deep] .271 .637 .822 .754 .214 .389 .792 .621 .865 18.8
±.${}^{\pm\text{.002}}$.${}^{\pm\text{.008}}$.${}^{\pm\text{.027}}$.${}^{\pm\text{.026}}$.${}^{\pm\text{.002}}$.${}^{\pm\text{.023}}$.${}^{\pm\text{.028}}$.${}^{\pm\text{.010}}$.${}^{\pm\text{.012}}$${}^{\pm\text{1.2}}$
@@
±.${}^{\pm\text{.002}}$.${}^{\pm\text{.006}}$.${}^{\pm\text{.005}}$.${}^{\pm\textbf{.001}}$.${}^{\pm\text{.041}}$.${}^{\pm\text{.003}}$.${}^{\pm\textbf{.004}}$.${}^{\pm\text{.003}}$.${}^{\pm\text{.002}}$${}^{\pm\text{1.2}}$
EVResNet50 .379 .808 .881 .913 .719 .635 .736 .839 .933 46.3
±.${}^{\pm\textbf{.002}}$.${}^{\pm\textbf{.001}}$.${}^{\pm\textbf{.002}}$.${}^{\pm\text{.004}}$.${}^{\pm\textbf{.005}}$.${}^{\pm\textbf{.007}}$.${}^{\pm\text{.009}}$.${}^{\pm\textbf{.003}}$.${}^{\pm\textbf{.012}}$${}^{\pm\textbf{1.2}}$
Among the suite of metrics used to assess alignment between models and primate visual behavior, shape bias has emerged as a particularly informative proxy of human-aligned inductive biases and out-of-domain (OOD) generalization [geirhos2022imagenettrained, geirhos2021partial]. Motivated by the hypothesis that shape bias may originate as early as early visual computations, we evaluated shape bias in EVNets (Tab. 1) and observed a substantial increase of 27.5% relative to the standard ResNet50 and of 14.7% relative to the VOneNet model, suggesting that EVNets may confer not only improved neuronal alignment but also behavioral traits more consistent with primate perception.
@@ 3.2 EVNets Improve Robustness Across an Aggregate Benchmark
When tested on clean images (Tab. 2), our VOneNets variant achieves a 1.2% improvement over the original VOneNet [Dapello2020.06.16.154542], while EVNets maintain the same reported accuracy. We further examined whether the increased alignment of EVNets with primate vision, also leads to improved robustness.
@@
Across most corruption categories and in terms of mean corruption accuracy, EVNets consistently outperformed both VOneNets and the base ResNet50, which only retained an advantage in weather corruptions. Notably, the most pronounced gains were observed on noise corruptions, where EVNets overperformed VOneNets by 2.5% while effectively preserving the same cumulative Fano factor as VOneNets.
Table 2: EVResNet50 overperforms baselines on most image corruption types and on mean corruption accuracy. Clean and corrupted top-1 accuracies averaged across severities and corruptions for ResNet50, VOneResNet50 and EVResNet50. Values indicate mean $\pm$ SD ($n=3$ seeds).
@@
Table 3 summarizes our OOD generalization results. EVNets consistently outperforms both the baseline ResNet50 and VOneNets across the majority of benchmarks, also surpassing these baselines on the mean domain shift accuracy. The sole exception is ImageNet-Sketch, where the base ResNet50 exhibits a marginal advantage.
[…]
@@
Given the lower clean accuracy of EVNets and prior findings that downstream stochasticity is the main contributor to VOneNet’s adversarial robustness [Dapello2020.06.16.154542], we anticipated a modest reduction in robustness relative to VOneNets. Interestingly, while VOneNets maintained the highest mean accuracy across all perturbation norms and attack strengths (Tab. 4), EVNets exhibited a smaller gap between clean and adversarial accuracy. Moreover, EVNets provided the greatest gains under stronger attacks, particularly within the $L_{2}$ and $L_{1}$ norm constraints, suggesting a potential link between training under inherited noise and robustness under larger perturbations.
Table 4: EVResNet50 overperforms base ResNet50 model under a range of adversarial perturbations and increases robustness to higher $L_{1}$ and $L_{2}$ norm constraints, when compared to VOneNets. Top-1 accuracies for the ResNet50, VOneResNet50 and EVResNet50 models. Values indicate mean $\pm$ SD ($n=2$ seeds).
@@
Mean $\frac{1}{1020}$ $\frac{1}{255}$ $\frac{4}{255}$ $\frac{16}{255}$ 0.15 0.6 1.2 2.4 40 160 640 2560
@@
ResNet50 [he2015deep] 8.5 24.1 0.6 0.2 0.2 37.6 2.1 0.2 0.2 34.2 1.9 0.2 0.2
±0.3 ±0.7 ±0.0 ±0.0 ±${}^{\pm{\textbf{0.0}}}$1.0 ±0.3 ±0.0 ±0.0 ±0.7 ±0.2 ±0.0 ±0.1
VOneResNet50 26.6 62.0 29.5 1.1 0.0 66.7 45.0 4.6 0.0 65.8 39.8 4.2 0.0
±${}^{\pm{\textbf{0.1}}}$${}^{\pm{\textbf{0.1}}}$${}^{\pm{\textbf{0.1}}}$${}^{\pm{\textbf{0.1}}}$0.0 ±${}^{\pm{\textbf{0.3}}}$${}^{\pm{\textbf{0.8}}}$0.1 ±0.0 ±${}^{\pm{\textbf{0.5}}}$${}^{\pm{\textbf{0.4}}}$0.1 ±0.0
EVResNet50 25.9 60.4 29.2 0.9 0.0 63.2 40.6 8.7 3.9 60.0 33.7 6.6 3.8
±0.0 ±0.3 ±0.0 ±0.1 ±0.0 ±0.3 ±0.4 ±${}^{\pm{\textbf{0.2}}}$${}^{\pm{\textbf{0.1}}}$0.1 ±0.4 ±${}^{\pm{\textbf{0.3}}}$${}^{\pm{\textbf{0.1}}}$
Table 5 summarizes these results and presents the Robustness Score for the evaluated models.
@@ 3.3 Combining EVNets with Data Augmentation Provides Cumulative Gains
In our evaluation of the Robustness Score for the EVResNet50 with PRIME data augmentation (Tab. 5), we observe the same performance gain as when PRIME is added to the standard ResNet50, exhibiting near-linear cumulative improvements. Furthermore, when comparing these results with those obtained by augmenting a ResNet50 with adversarial training, no configuration surpasses the additive gains achieved by the EVResNet50 trained with PRIME.
@@
ResNet50 [he2015deep] 27.0$\pm$0.1 8.6$\pm$0.3 39.1$\pm$0.3 33.3$\pm$0.1 75.4$\pm$0.1
VOneResNet 34.6$\pm$0.1 26.6$\pm$0.1 40.4$\pm$0.2 36.9$\pm$0.1 72.9$\pm$0.0
EVResNet 35.4$\pm$0.0 25.9$\pm$0.0 41.5$\pm$0.1 38.7$\pm$0.1 71.7$\pm$0.1
[…]
ResNet50 + PRIME [10.1007/978-3-031-19806-9_36] 34.2 14.8 52.6 42.6 76.0
EVResNet50 + PRIME 41.5$\pm$0.2 27.4$\pm$0.2 52.6$\pm$0.0 44.7$\pm$0.3 73.0$\pm$0.1
3.4 Not All Components of the SubcorticalBlock Contribute to Improving Robustness
Though ablating components of the SubcorticalBlock in most cases reduced model robustness (cf. Supplementary Material D.1), removing the M cell pathway had minimal effect.
4 Discussion
A central challenge in computer vision lies in designing models that are robust to diverse image perturbations. This is not only an essential property for real-world deployment but also for approximating biological vision. In this work, we present a new family of neuro-inspired CNNs that not only display enhanced robustness across a broad spectrum of perturbations but also achieve stronger alignment with primate vision. In contrast to our approach, the current literature has relied on combining CNN architecture with isolated components of early visual processing, such as RF filters [10.1007/978-3-031-44204-9_33, EVANS202296] and normalization schemes [NIPS2012_c399862d, 5459469]. Our SubcorticalBlock, however, draws directly from computational neuroscience, with parameters constrained by decades of empirical findings. The EVNet architecture introduces a modular, cascading model of V1 that incorporates stage-specific architectural priors, offering a compelling alternative to V1-inspired CNN paradigms. While EVNets do not entirely solve the longstanding problem of robust generalization, our results underscore the critical importance of subcortical processing in shaping early visual representations. Notably, we demonstrate that meaningful improvements to V1 modeling can be achieved by exclusively refining upstream stages. The cumulative gain shown by combining the SubcorticalBlock with the VOneBlock reveal a biologically-plausible potential for compositionality in inductive biases, where each module targets distinct axes of visual invariance, together producing synergistic improvements in OOD generalization. Specifically, while the SubcorticalBlock primarily encodes invariance to luminance and contrast, the VOneBlock focuses more on spatial and polarity invariance. Finally, we show that integrating biologically-inspired architectures with standard data augmentation techniques leads to synergistic improvements in robustness. This suggests that neuroscience-driven inductive biases and machine learning heuristics are not mutually exclusive, but can in fact be complementary.
While the observed robustness gains are compelling, there are also some trade-offs. Specifically, our model abstracts subcortical processing by instantiating only four channels that reflect the average spatial response profiles observed in the LGN, rather than attempting to capture the full heterogeneity of subcortical cells. A related limitation of our approach is that M-cell pathway contributes only marginally to downstream performance (cf. Supplementary Material D.1), which, while being in line with the classical view of M cells contribution to the ventral stream, also presents inconsistencies with current literature [CANARIO2016110]. Additionally, enhancements in perturbation robustness are consistently accompanied by modest reductions in clean image accuracy — a common tension in robustness research [pushpull2024, Dapello2020.06.16.154542]. Moreover, the slight decrease in adversarial robustness may point to a deeper dissociation between inherited and V1-intrinsic noise in the VOneBlock architecture, opening a promising avenue for future investigation. It is important to mention that the noise injected at the SubcorticalBlock has considerably lower dimensionality than that in V1 (4 channels vs. 512 channels, across the spatial map). Beyond these considerations, subsequent work could examine the generalization of these robustness benefits across diverse training distributions. In terms of V1 alignment, it is important to note that no changes were made in the VOneBlock to accommodate upstream processing. We opted for this approach to facilitate model comparison and because there were no drastic changes to VOneBlock responses other than the improvement of extra-classical RF phenomena. However, alignment to V1 decreased in a couple of benchmarks which is likely due to the interaction of the two modules. Future work, could explore how the GFB in the VOneBlock should change to reflect subcortical preprocessing. Finally, as EVNets modulate only subcortical extra-classical phenomena, it remains an open question whether incorporating V1-inspired normalization mechanisms could further enhance V1 alignment.
Broader impact.
This work advances computer vision by providing a biologically-plausible alternative to the increasingly popular, but non-biologically-based, transformer architectures [NIPS2017_3f5ee243, dosovitskiy2021an]. By simulating early visual processes, we narrow the gap between human and machine vision, fostering greater robustness while enhancing transparency and interpretability, besides a potential to reduce bias and enhance accessibility in underrepresented settings. Nevertheless, fully replicating human vision remains a challenge, requiring further research and more elaborate modeling. It is important to acknowledge that some of the foundational studies informing our study involved animal research. While these studies have contributed vastly to our understanding of the visual system, they also highlight the ethical considerations. In line with contemporary ethical standards, we recognize and underscore the importance of minimizing harm and using alternative methods whenever possible. This work illustrates how leveraging existing biological data can inform model development, potentially reducing reliance on new animal experiments.
Acknowledgments
This work was supported by the project Center for Responsible AI reference no. C628696807-00454142, financed by the Recovery and Resilience Facility, by project PRELUNA, grant PIDC/CCIINF/4703/2021, and by Fundação para a Ciência e Tecnologia (FCT), under grant UIDB/50021/2020.

### v2

@@ 3.1 EVNets Improve Neuronal and Behavioral Alignment with Primate Vision
We evaluated whether coupling the SubcorticalBlock with the VOneBlock improved alignment with primate V1. As illustrated in Figure 2, the inclusion of the SubcorticalBlock upstream of the VOneBlock introduces hallmark extra-classical RF response properties absent from the VOneBlock alone. In particular, we observe an increased surround suppression in the size tuning curve and a non-linear contrast-sensitivity curve.
Figure 2: Subcortical preprocessing improves explanability of extra-classical RF properties in V1. SF, size, and contrast tuning curves (left to right) for an example neuron in the VOneBlock with and without subcortical preprocessing (M cell). Example frames from the drifting gratings stimuli are shown below. VOneBlock neurons in isolation exhibit predominantly classical RF effects but when coupled with subcortical processing exhibit behaviors consistent with those empirically observed, such as enhanced surround modulation and non-linear contrast responses[66, 67]. See Supplementary Material B.1 for empirical V1 tuning curves.
Motivated by these observations, we used the BrainScore platform [42, 43] to quantitatively evaluate V1 alignment. As shown in Table 1, while VOneNets outperform in V1 predictivity, EVResNet50 achieves a higher mean response property score than both the ResNet50 and VOneResNet50 models. Notably, the highest gains are observed for the surround modulation and RF size tuning, both associated with the increased surround suppression. A tradeoff in SF tuning is observed, which can be attributed to the fact that no changes were done to the GFB to account for the upstream processing.
Table 1: EVResNet50 outperforms baselines on mean V1 response property alignment, and shape bias. BrainScore [42, 43] V1 alignment scores and shape bias [8] for ResNet50, VOneResNet50, and EVResNet50. Values indicate mean $\pm$ SD ($n=3$ seeds).
@@
ResNet50 .271 .637 .822 .754 .214 .389 .792 .621 .865 18.8
$^{\pm\text{.002}}$ $^{\pm\text{.008}}$ $^{\pm\text{.027}}$ $^{\pm\text{.026}}$ $^{\pm\text{.002}}$ $^{\pm\text{.023}}$ $^{\pm\text{.028}}$ $^{\pm\text{.010}}$ $^{\pm\text{.012}}$ $^{\pm\text{1.2}}$
@@
$^{\pm\text{.002}}$ $^{\pm\text{.006}}$ $^{\pm\textbf{.005}}$ $^{\pm\textbf{.001}}$ $^{\pm\text{.041}}$ $^{\pm\text{.003}}$ $^{\pm\textbf{.004}}$ $^{\pm\textbf{.003}}$ $^{\pm\text{.002}}$ $^{\pm\text{1.2}}$
EVResNet50 .364 .826 .854 .950 .726 .614 .916 .781 .933 48.9
$^{\pm\text{.000}}$ $^{\pm\textbf{.000}}$ $^{\pm\text{.009}}$ $^{\pm\text{.000}}$ $^{\pm\textbf{.001}}$ $^{\pm\textbf{.004}}$ $^{\pm\text{.001}}$ $^{\pm\text{.000}}$ $^{\pm\textbf{.000}}$ $^{\pm\textbf{2.4}}$
Among the suit of primate visual behavior alignment metrics, shape bias has emerged as a particularly informative proxy of human-aligned inductive biases and out-of-domain (OOD) generalization [8, 7]. Motivated by the hypothesis that shape bias may originate in early visual computations, we evaluated shape bias in EVNets (Tab. 1) and observed a substantial increase of 30.1% relative to the standard ResNet50 and of 17.3% relative to the VOneNet model, suggesting that EVNets may confer not only improved neuronal alignment but also behavioral traits more consistent with primate perception.
@@ 3.2 EVNets Improve Robustness Across an Aggregate Benchmark
When tested on clean images (Tab. 2), our VOneNets variant achieves a 1.2% improvement over the original VOneNet [12], while EVNets displayed a performance drop of 1.3%, when compared to the same reported accuracy. We further examined whether the increased alignment of EVNets with primate vision, also leads to improved robustness.
@@
Across most corruption categories and in terms of mean corruption accuracy, EVNets consistently outperformed both VOneNets and the base ResNet50, which only retained an advantage in weather corruptions. Notably, the most pronounced gains were observed on noise corruptions, where EVNets outperformed VOneNets by 3.7% while effectively preserving the same cumulative Fano factor as VOneNets (cf. Fig. C4 for accuracy across individual corruptions).
Table 2: EVResNet50 outperforms baselines on most image corruption types and on mean corruption accuracy. Clean and corrupted top-1 accuracies averaged across severities and corruptions for ResNet50, VOneResNet50 and EVResNet50. Values indicate mean $\pm$ SD ($n=3$ seeds).
@@
Table 3 summarizes our OOD generalization results. EVNets consistently outperforms both the baseline ResNet50 and VOneNets across the majority of benchmarks, also surpassing these baselines on the mean domain shift accuracy. The only dataset where EVNets underperform is ImageNet-Sketch, where the base ResNet50 exhibits a marginal advantage.
Table 3: EVResNet50 outperforms baselines on most domain shift datasets and on overall mean domain shift accuracy. Top-1 accuracies on ImageNet-{Cartoon, Drawing, R, Sketch, Stylized16} for ResNet50, VOneResNet50 and EVResNet50. Values indicate mean $\pm$ SD ($n=3$ seeds).
@@
EVNets improves adversarial robustness across most perturbation norms and attack strengths when compared to VOneNets (Tab. 4). While VOneNets obtained a marginal advantage under the weakest attack strengths for $L_{\infty}$ and $L_{2}$ norm constraints, EVNets improved robustness on the remaining attack settings, exhibiting also a smaller gap between clean and adversarial accuracy.
Table 4: EVResNet50 outperforms baselines on most adversarial perturbations and on mean adversarial robustness. Top-1 accuracies for the ResNet50, VOneResNet50 and EVResNet50 models. Values indicate mean $\pm$ SD ($n=3$ seeds).
@@
Mean $\frac{1}{1020}$ $\frac{1}{255}$ $\frac{4}{255}$ $\frac{16}{255}$ 0.15 0.6 2.4 9.6 40 160 640 2560
@@
ResNet50 8.3 23.4 0.4 0.2 0.2 37.2 1.8 0.2 0.2 33.6 1.7 0.2 0.2
±0.2 ±0.8 ±0.0 ±0.0 $^{\boldsymbol{\pm}\textbf{0.0}}$ ±1.0 ±0.2 ±0.0 $^{\boldsymbol{\pm}\textbf{0.0}}$ ±0.7 ±0.2 ±0.0 $^{\boldsymbol{\pm}\textbf{0.0}}$
VOneResNet50 26.1 62.6 30.4 1.2 0.0 66.2 42.3 4.2 0.0 64.5 37.3 3.9 0.0
$^{\boldsymbol{\pm}\textbf{0.1}}$ ±0.4 ±0.3 ±0.1 ±0.0 $^{\boldsymbol{\pm}\textbf{0.3}}$ ±0.2 ±0.1 ±0.0 $^{\boldsymbol{\pm}\textbf{0.3}}$ ±0.6 ±0.2 ±0.0
EVResNet50 28.3 62.7 38.8 3.0 0.0 65.1 48.0 7.4 0.0 64.0 44.5 6.0 0.0
$^{\boldsymbol{\pm}\textbf{0.2}}$ $^{\boldsymbol{\pm}\textbf{0.2}}$ $^{\boldsymbol{\pm}\textbf{0.6}}$ $^{\boldsymbol{\pm}\textbf{0.1}}$ ±0.0 ±0.0 $^{\boldsymbol{\pm}\textbf{0.3}}$ $^{\boldsymbol{\pm}\textbf{1.8}}$ ±0.0 ±0.2 $^{\pm\textbf{0.4}}$ $^{\pm\textbf{1.8}}$ ±0.0
Table 5 summarizes all the robustness results described above and presents the Robustness Score for the evaluated models. The EVResNet50 model improves the Robustness Score by 9.3% over the base ResNet50 and by 1.6% over our VOneNet variant.
@@ 3.3 Combining EVNets with Data Augmentation Provides Cumulative Gains
In our evaluation of the Robustness Score for the EVResNet50 with PRIME data augmentation (Tab. 5), we find that this combined strategy yields cumulative performance gains beyond those of either component alone. Furthermore, when comparing these results with those obtained by augmenting a ResNet50 with adversarial training and augmenting a VOneNet with PRIME, no configuration surpasses the additive gains achieved by the EVResNet50 trained with PRIME.
@@
[…]
3.4 EVNet Variants Reveal Competing Drivers of Primate Vision Alignment and Robustness
Selective ablation of specific SubcorticalBlock components affected primate vision alignment in different aspects (see Supplementary Material D.1). The components that caused a greater drop in V1 response property benchmarks were the M-cell pathway and the contrast normalization, which greatly affected the RF size and surround modulation. Surprisingly, removing either the P-cell pathway or light adaptation greatly increased shape bias.
Table 6: Model components contribution to robustness vary greatly. Robustness Score, clean and perturbed top-1 accuracies for all EVResNet50 variants. ResNet50 and VOneResNet50 included for reference but not in the comparison. Values indicate mean $\pm$ SD ($n=2$ seeds).
Perturbations
Robust. Score Adversarial Corrupt. Domain Shift Clean
Model [%] [%] [%] [%] [%]
[…]
Computed with a reduced attack set (two perturbations per norm constraint); not comparable to Table 5.
In terms of performance, the single component that had the largest impact on both clean accuracy and robustness was the P-cell pathway (Tab. 6). On the other hand, removing the M-cell pathway or contrast normalization had no impact on clean accuracy and robustness. Light adaptation had only minor effects on specific robustness benchmarks and a small drop in clean accuracy. Removing subcortical stochasticity improved clean accuracy while decreasing adversarial robustness. Interestingly, the omission of the VOneBlock barely affected overall robustness, revealing even to be the best variant under image corruptions and an improvement in clean accuracy, accompanied by a small drop in adversarial robustness. The inclusion of LGN-V2 skip connections had little to no impact in performance across all perturbations types and clean images.
4 Discussion
In this work, we present a new family of neuro-inspired CNNs that not only display enhanced robustness across a broad spectrum of perturbations but also achieve stronger alignment with primate vision. Previous works have incorporated biologically inspired mechanisms such as DoG filtering [30, 16, 15], divisive normalization [13, 68], and noise injection [37] within CNN architectures. In contrast, our SubcorticalBlock introduces a principled segregation of computations into parallel P- and M-cell pathways constrained by prior empirical neurophysiological findings. The EVNet architecture introduces a modular, cascading model of V1 that incorporates stage-specific architectural priors, offering a compelling alternative to V1-inspired CNN paradigms. The V1-alignment improvements observed in EVNets are most pronounced for surround modulation and RF size. These effects likely arise from the normalization mechanisms within the SubcorticalBlock, which induce local competition and lead to extraclassical RF responses not present in VOneNets. Despite these improvements, benchmark scores for surround modulation and RF size are still relatively low, suggesting that normalization mechanisms at the V1 level, either caused by recurrent or feedback circuits, are also needed for a better alignment [13, 68]. While EVNets do not entirely solve the longstanding problem of robust generalization, our results also underscore the critical importance of subcortical processing in shaping early visual representations: while the DoG filtering improves features selectivity, performs low-pass filtering, mitigating high-frequency noise, the normalization layers promote local competition and dynamic range compression, reducing sensitivity to input perturbations. Notably, we demonstrate that meaningful improvements to V1 modeling can be achieved by exclusively refining upstream stages. The cumulative gain shown by combining the SubcorticalBlock with the VOneBlock reveals a biologically-plausible potential for compositionality in inductive biases, where each module targets distinct axes of visual invariance, together producing synergistic improvements in perturbation robustness. Specifically, while the SubcorticalBlock primarily encodes invariance to luminance and contrast, the VOneBlock focuses more on spatial and polarity invariance. Finally, we show that integrating biologically-inspired architectures with standard data augmentation techniques leads to synergistic improvements in robustness, surpassing data augmentation alone and adversarial training, the most effective methods for improving adversarial and corruption robustness. Together, these results provide further evidence that neuroscience-driven inductive biases and machine learning heuristics are not mutually exclusive, but can in fact be complementary.
While the observed robustness gains are compelling, there are also some trade-offs. Specifically, our model abstracts subcortical processing by instantiating only four channels that reflect the average spatial response profiles observed in the LGN, rather than capturing the full heterogeneity of subcortical cells. An interesting observation is that the M-cell pathway contributes marginally to downstream performance, which is in line with the classical view of M cells small contribution to the ventral stream. However, this result presents some inconsistencies with more recent literature [69]. Additionally, enhancements in perturbation robustness are consistently accompanied by modest reductions in clean image accuracy, a common tension in robustness research [14, 12]. Beyond these considerations, subsequent work could examine the possibility of initializing from neuro-inspired weights while allowing task-driven fine-tuning of the front-ends used. For V1 alignment, we kept the VOneBlock unchanged as upstream processing minimally affected its responses aside from the improved extra-classical RF effects. However, the slight drop in V1 alignment on some benchmarks likely reflects unforseen interactions between modules. Future work could adapt the GFB to account for subcortical preprocessing and add normalization mechanisms at the V1 level to evaluate if these changes further enhance alignment and robustness. Finally, while our approach may seem to contrast the “Bitter Lesson” [70], the trajectory of neurally aligned vision tells a more nuanced story. Prior work indicates that scaling alone does not yield more brain-like representations [71], and large-scale benchmarks [72] show that neural alignment with primate visual cortex saturates with model size — architectures with stronger biological priors often align better than transformers. Our work leverages this insight by embedding early vision–inspired inductive biases to achieve the efficiency and alignment characteristic of biological vision.
Broader impact.
This work advances computer vision by proposing a biologically grounded alternative to vision transformers [73]. By simulating early vision, we bridge the gap between human and machine vision, improving robustness, transparency, and interpretability, with potential benefits for bias reduction and accessibility. Fully replicating human vision remains challenging and calls for further research. We acknowledge that some foundational insights stem from animal studies, emphasizing the ethical responsibility to minimize harm and favor alternative approaches. Leveraging existing biological data, as done here, may reduce reliance on new animal experimentation.
Acknowledgments
This work was supported by the project Center for Responsible AI reference no. C628696807-00454142, and by national funds through Fundação para a Ciência e a Tecnologia, I.P. (FCT) under projects UID/50021/2025 and UID/PRR/50021/2025.